Q. In air, a metallic sphere with an internal cavity weighs 40 g and in water it weighs 20 g. What is the volume of the cavity if the density of material with cavity is 8 g/cm3

Grade:Select Grade

1 Answers

Sai Sandeep
9 years ago
First calculate volume of solid part for weight in air and denisty of material.
Vs=5 cc.When it is placed in water net pward force=net downward force
=>40g-(V+Vs)g=20g =>V+Vs=20 therefore V=20-Vs=15